# LDMol: Text-to-Molecule Diffusion Model with Structurally Informative Latent Space

### v2

@@ A.3 More detailed procedure for LDMol’s application on downstream tasks
Algorithm 1 Molecule-to-Text Retrieval with LDMol
1:$z,\mathcal{C}=\{c_{i}\}^{B}_{i=1},n\in\mathbb{N}^{+}$
2:Initialize $\text{Initialize }\texttt{Errors}[c_{i}]=0\textbf{ for }i=1\textbf{ to }B$
3:for iter=1 to $\texttt{iter}=1\textbf{ to }n$ do
4:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
5:     $z_{t}=\sqrt{\overline{\alpha}_{t}}z+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
6:     for $i=1\textbf{ to }B$ do
7:         $\texttt{Errors}[c_{i}]\mathrel{+}=||\hat{\epsilon}_{\theta}(z_{t},t,c_{i})-\epsilon||_{2}^{2}$
8:     end for
9:end for
10:Return $\text{argmin}_{c_{i}\in\mathcal{C}}\texttt{Errors}[c_{i}]$
Algorithm 2 Text-guided Molecule Editing with LDMol
1:$z_{src},c_{src},c_{tgt},N\in\mathbb{N}^{+},\gamma>0,\omega\geq 1,\mathcal{D}$
2:Initialize $\text{Initialize }z_{tgt}=z_{src}$
3:for iter=1 to $\texttt{iter}=1\textbf{ to }N$ do
4:     $t\sim U[0,T],\epsilon\sim\mathcal{N}(0,I)$
5:     $z_{t,src}=\sqrt{\overline{\alpha}_{t}}z_{src}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
6:     $z_{t,tgt}=\sqrt{\overline{\alpha}_{t}}z_{tgt}+\sqrt{1-\overline{\alpha}_{t}}\epsilon$
7:     $\epsilon_{\theta,src}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,src},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,src},t,c_{src})$
8:     $\epsilon_{\theta,tgt}^{\omega}=(1-\omega)\epsilon_{\theta}(z_{t,tgt},t,\varnothing)+\omega\epsilon_{\theta}(z_{t,tgt},t,c_{tgt})$
9:     $z_{tgt}=z_{tgt}-\gamma(\epsilon_{\theta,tgt}^{\omega}-\epsilon_{\theta,src}^{\omega})$
10:end for
11:Return $\mathcal{D}(z_{tgt})$
@@
Figure 7-(b) and Algorithm 2 illustrate the DDS-based molecule editing with LDMol. Specifically, DDS requires source data $z_{src}$, target data $z_{tgt}$ which is initialized to $z_{src}$, and their corresponding source and target text descriptions $\{c_{src},c_{tgt}\}$. We apply the forward noise process to $z_{src}$ and $z_{tgt}$ using the same randomly sampled $t$ and $\epsilon$ to get $z_{t,src}$ and $z_{t,tgt}$. These are fed into the pre-trained LDMol with their corresponding text, where we denote the output noise as $\hat{\epsilon}_{\theta,src}$ and $\hat{\epsilon}_{\theta,tgt}$. Finally, $z_{tgt}$ is modified towards the target text by optimizing it to the direction of $(\hat{\epsilon}_{\theta,tgt}-\hat{\epsilon}_{\theta,src})$ with a learning rate $\gamma$. Here, $\hat{\epsilon}_{\theta,tgt}$ and $\hat{\epsilon}_{\theta,tgt}$ can be replaced with the classifier-free-guidanced noises, utilizing the output with the null text and the guidance scale $\omega$. $z_{tgt}$ is decoded back as the editing output after the optimization step is iterated $N$ times. In Figure 5-(a), where we applied the same scenario to a batch of molecules, we used null text as $c_{src}$ since it’s impractical to prepare a source prompt for each molecule. The hyperparameters $\{N,\gamma,\omega\}$ are fixed for each scenario, where every choice is in the range of $100\leq N\leq 200$, $0.1\leq\gamma\leq 0.3$ and $2.0\leq\omega\leq 4.5$. For the illustrated examples in Figure 5-(b), we repeated DDS iterations with $N=150$, $\gamma=0.1$ and $\omega=2.5$.
@@ Appendix B Additional results
B.1 Effect of classifier-free guidance scale on LDMol sample quality
Figure 8 plots the LDMol’s text-to-molecule generation performance on the ChEBI-20 test set, with different classifier-free guidance scale $\omega$ in the sampling process. Starting from $\omega=1.0$ which is equivalent to a naive conditional generation, we observed the overall sample quality is improved as $\omega$ increases but collapses for too big $\omega$. This agrees with the well-known observation on image diffusion models, and we decided to use $\omega=2.5$ for the text-to-molecule generation with LDMol.
@@ B.1 Effect of classifier-free guidance scale on LDMol sample quality
B.2 Molecule-to-text retrieval examples

### v4

@@ A.3 LDMol’s application on downstream tasks
Figure 8-(b) and Algorithm 2 illustrate the DDS-based molecule editing with LDMol. Specifically, DDS requires source data $z_{src}$, target data $z_{tgt}$ which is initialized to $z_{src}$, and their corresponding source and target text descriptions $\{c_{src},c_{tgt}\}$. We apply the forward noise process to $z_{src}$ and $z_{tgt}$ using the same randomly sampled $t$ and $\epsilon$ to get $z_{t,src}$ and $z_{t,tgt}$. These are fed into the pre-trained LDMol with their corresponding text, where we denote the output noise as $\hat{\epsilon}_{\theta,src}$ and $\hat{\epsilon}_{\theta,tgt}$. Finally, $z_{tgt}$ is modified towards the target text by optimizing it to the direction of $(\hat{\epsilon}_{\theta,tgt}-\hat{\epsilon}_{\theta,src})$ with a learning rate $\gamma$. Here, $\hat{\epsilon}_{\theta,tgt}$ and $\hat{\epsilon}_{\theta,tgt}$ can be replaced with the classifier-free-guided noises, utilizing the output with the null text and the guidance scale $\omega$. $z_{tgt}$ is decoded back as the editing output after the optimization step is iterated $N$ times. In Figure 5, where we applied the same scenario to a batch of molecules, we used null text as $c_{src}$ since it’s impractical to prepare a source prompt for each molecule. The hyperparameters $\{N,\gamma,\omega\}$ are fixed for each scenario, where every choice is in the range of $100\leq N\leq 200$, $0.1\leq\gamma\leq 0.3$ and $2.0\leq\omega\leq 4.5$. Following MoleculeSTM, each scenario was applied to 200 randomly sampled molecules from ZINC15, and the mean and standard deviation on three separate runs were plotted.
@@ Appendix B Additional results
B.1 Visualization of the LDMol latent space
Figure 9: UMAP visualization of LDMol encoder output (a) and the final latent space after the linear compression layer (b), from 10 groups containing 100 molecules each with shared Murcko scaffold(colored) and 5,000 general molecules(light grey).
To visualize the structural information encoded in the latent space of our encoder, we prepared 10 molecular clusters that contain 100 molecules, each sharing the common Murcko scaffold (Bemis & Murcko, 1996). Then, we obtained their latent vector from the LDMol encoder and visualized them in 2D via UMAP (McInnes et al., 2018). Note that As shown in Figure 9, the molecules with shared Murcko scaffold have formed clusters in the latent vector space.
B.2 Effect of classifier-free guidance scale
Figure 10 plots the LDMol’s text-to-molecule generation performance on the ChEBI-20 test set, with different classifier-free guidance scale $\omega$ in the sampling process. Starting from $\omega=1.0$, which is equivalent to a naive conditional generation, we observed that the overall sample quality is improved as $\omega$ increases but collapses for too big $\omega$. This agrees with the well-known observation on image diffusion models, and we decided to use $\omega=2.5$ for the text-to-molecule generation with LDMol.
@@ B.2 Effect of classifier-free guidance scale
B.3 Text-to-molecule generation
We measured the uniqueness, novelty, and prompt alignment score for the prompts in Figure 4 using 1,000 samples. Validity is the proportion of generated SMILES that are valid. Uniqueness is the proportion of valid SMILES that are unique. The “align” score is the proportion of unique SMILES that match the given prompt. Novelty is the proportion of the unique SMILES that are not included in the training dataset. The alignment score was measured with SMILES pattern matching with the substructure described by the prompt. We observed that even when stochastic sampling was enabled, AR models struggled to generate various samples from a single prompt. LDMol can generate molecules that align better with various hand-written prompts. Furthermore, its outputs were much more diverse than the previous AR models.
Figure 12 shows the behavior of LDMol’s text-to-molecule generation with several exceptional scenarios. When we fed a completely ambiguous input such as “beautiful” or “important”, the model spits out a variety of different molecules without any consistency. When we fed contradictory inputs that could not be satisfied, the outputs were chimeric between contradictory prompts, with a clearly decreased validity.
Table 6: Quantitative results of the case studies in Figure 4. The best performance for each metric is written in bold.
Models Validity(V) Uniqueness(U) Align(A) V×$\times$×U×$\times$×A Novelty
Case (a) molT5large 0.996 0.006 1.000 0.006 -
bioT5+ 0.846 0.028 0.625 0.015 -
LDMol 0.910 0.951 1.000 0.865 0.988
Case (b) molT5large 0.927 0.012 0.818 0.009 -
bioT5+ 1.000 0.573 0.782 0.448 -
LDMol 0.989 0.960 0.906 0.860 0.958
Case (c) molT5large 0.783 0.072 0.643 0.036 -
bioT5+ 1.000 0.160 0.750 0.120 -
LDMol 0.955 0.861 0.688 0.566 0.780
Case (d) molT5large 0.995 0.002 0.500 0.001 -
bioT5+ 1.000 0.015 0.733 0.011 -
LDMol 0.956 0.849 0.703 0.571 0.842
Case (e) molT5large 0.956 0.015 0.571 0.008 -
bioT5+ 1.000 0.035 0.086 0.003 -
LDMol 0.996 0.187 0.595 0.111 0.667
Figure 11: Examples of the generated molecules by LDMol, with (a, b) ambiguous text conditions and (c) contradictory and unreasonable input.
Figure 12: Examples of text-guided molecule editing with LDMol. The difference between the source text and the target text, and the corresponding region, is colored in purple.
B.4 Molecule-to-text retrieval
@@
B.5 Text-guided molecule editing
Figure 12 illustrates several case studies with hand-written editing prompts and results, where the editing output successfully modified the input molecule towards the target prompt with minimal corruption of the unrelated region. Here, we repeated DDS iterations with $N=150$, $\gamma=0.1$ and $\omega=2.5$.
B.6 Ablation study
Table 7: Quantitative results of the ablation study. The best performance for each metric is written in bold.
models Autoencoder ChEBI-20 text-to-molecule generation
[…]
LDMol w/o compression layer 0.964 0.022 0.000 67.93
LDMol w/ transformer compression layer 0.986 0.565 0.084 2.19
LDMol w/o stereoisomer hard-negative 0.891 0.939 0.278 0.24
LDMol 0.983 0.941 0.530 0.20
We’ve conducted an ablation study on more detailed design choices of the proposed LDMol in Table 7 to analyze and emphasize their role.
When we didn’t introduce a compression layer, the later diffusion model completely failed to learn the latent space since its dimension was too big for the diffusion model to learn. We tried to utilize a more complex compression module by transformer encoder layers of Perceiver-Resampler (Alayrac et al., 2022; Lovelace et al., 2024) manner, but the performance was significantly decreased as shown in the second row. This is presumably because adding another complicated layer makes the latent space deviate from the former informative and well-regulated learnable space.
When stereoisomers were not utilized as hard negative samples in the contrastive encoder training, the constructed latent space was not detailed enough to specify the input, which degraded the reconstruction accuracy of the autoencoder. The similarity metric of FCD didn’t decrease as much, but the exact match ratio has decreased significantly.
B.7 Computational efficiency
Table 8: Quantitative results of the ablation study.
Models molT5large bioT5+ LDMol
Required time[s] 523 180 361
VRAM usage[GB] 4.92 1.08 3.79
Table 8 compares the computational efficiency of LDMol and several baselines with state-of-the-art performance. In terms of memory usage, our model can operate with less than 4GB of VRAM, which is smaller than that of molT5large. The required time was also comparable to transformer-based models, even with the latent decoder and the Classifier-Free Guidance(CFG) which doubles the diffusion model usage. Considering many works have been published to reduce the inference time of diffusion models (Song et al., 2023; Salimans & Ho, 2022), as one of the first successful text-to-molecule diffusion models, we believe that the inference time can be further improved by future research.